10.5, 14.6, and 18.2 Ohm resistors are connected in series to a 12.0V battery (you can ignore the internal resistance of the battery).. What is the current in Amps through the battery? (Assume 3 significant digits).
Answer #1

When resistances are connected in series ,the equivalent resistance is given as

R_{eq}=R_{1}+R_{2}+R_{3}=10.5+14.6+18.2=43.3\Omega

The voltage supplied is V=12 Volts

Thus, using Ohm's Law, Current through the battery = V/Req =12/43.3 =0.277 A